OpenStudy (anonymous):

simplify the expression. cos(theta)/1+sin(theta) + tan(theta)

OpenStudy (john_es):

You mean, \[\frac{\cos\theta}{1+\sin\theta+\tan\theta}\]?

OpenStudy (john_es):

If the expression was, \[\frac{\cos\theta}{1+\sin\theta}+\tan\theta\] it could be simplified, but the other expression I cannot find the way to simplify it.